# Large Language Models:

### v1

@@ 5.4.2 Monte Carlo Simulations based on Congressional Legislation
Figure 7 and the top panel of Table 4(b) compares the average bias of the plug-in coefficient $\widehat{\beta}$ and the bias corrected coefficient $\widehat{\beta}^{debiased}$ for the target regression $\beta^{*}$ (normalized by their standard deviations) across possible combinations of bill topic $V_{r}$, linked covariate $W_{r}$, LLM $m$, and prompting strategy $p$. For most regression specifications and pairs of LLM and prompting strategy, the simple plug-in regression suffers from substantial biases for the target regression. Indeed, focusing on GPT-3.5-Turbo, the 5-th percentile yields a bias that is fully $-1.86$ standard deviations away from the target regression coefficient and the 95th percentile is fully $2.18$ standard deviations off. By contrast, using the validation sample to debias the LLM labels yields estimates that are reliably unbiased for the target regression — indeed, the bias-corrected regression coefficient performs remarkably well across all regression specifications and pairs of LLM and prompting strategy.
We have shown above that measurement error in the LLM labels leads not only to biased point estimates, but correcting using a validation sample fixes this problem. Does this extend to our inference statements? For each regression specification and pair of LLM and prompting strategy, we further calculate the fraction of simulations in which a 95% confidence interval centered at either the plug-in coefficient or the bias-corrected coefficient includes the target regression $\beta^{*}$. The bottom panel of Table 4(b) summarizes the distribution of coverage results across all combinations of regression specification and pair of LLM and prompting strategy. For GPT-3.5-Turbo (left panel, bottom), the median $95\%$ confidence interval obtained from just plugging in the language model’s label includes the true parameter just $81\%$ of the time; the $5^{th}$ percentile confidence interval includes the true parameter just $38\%$ of the time. In contrast the $95\%$ confidence interval from the de-biased label covers the true parameter value around $95\%$ of the time the vast majority of the time. The right-hand panel shows qualitatively similar results with GPT-4o as well.
We can use these data to also answer the question: If we have already gone to the trouble of collecting some gold standard labels, what is the additional value of the LLM labels in the primary sample? We could now in principle just directly run the target regression on the validation sample.
@@
Figure 8 illustrates what could be gained by using the LLM on the primary sample. For each regression specification and pair of LLM and prompting strategy, we calculate the mean square error of the bias-corrected coefficient and the validation-sample only coefficient for the target regression $\beta^{*}$. We then calculate the average mean square error across all 1000 simulations for each regression specification and pair of LLM and prompting strategy, and Figure 8 plots the resulting distribution across all choices of bill topic $V_{r}$, covariate $W_{r}$, and pair of model-and-prompting strategy.
In this empirical setting, we find that the average mean square error of the validation-sample-only coefficient is always higher (less precisely estimated) compared to the bias-corrected coefficient. This suggests that there is information extracted from the LLM’s labels on the primary sample, and further illustrates that precision improvements, as predicted by the asymptotic comparison in Equation (14), are possible in finite samples. In Appendix D, we show that these precision gains dissipate once the validation sample is around 25% of the researcher’s sample. The share of applications in which it would actually be feasible (from a time and cost perspective) to collect gold-standard labels for more than a quarter of the sample is an open question. We suspect that for many applications, a combination of LLM labels and a small amount of validation data is likely to be the optimal approach.
These results focused on using the LLM’s labels for the economic concept $V_{r}$ as the dependent variable in a linear regression. As we show in Appendix D, we find similar results for the case in which the economic concept is an explanatory variable in a regression. In particular, we continue to find that plug-in regression with LLM labels leads to substantial biases and coverage distortions; by contrast, bias-correcting using a small validation sample effectively eliminates bias and restores the coverage of conventional confidence intervals. For many regression specifications and pair of LLM and prompting strategy, we again find that the bias-corrected coefficient improves on the mean-square error of the validation-sample-only coefficient.
Taken together, these results highlight that LLM outputs are not substitutes for gold standard labels. But rather, when used correctly, LLM outputs can amplify a small validation sample, enabling the researcher to draw correct and precise inferences at a lower cost.
6 Extensions to Novel Uses of LLMs
So far we have applied our framework to familiar empirical problems related to prediction and estimation. But LLMs are increasingly being used in ways entirely new to economics research, such as for automated hypothesis generation, or to simulate the responses of human subjects. In this section, we show that our framework is broad enough to cover even novel LLM uses like these, partly because these new uses can often be viewed as variants of prediction and estimation problems.
6.1 LLMs as Hypothesis Generators
In Section 4, we considered whether LLM outputs can be used for prediction — the task at the heart of supervised learning. By reinterpreting its components, our framework also gives us a way to understand recent work using LLMs as tools for hypothesis generation. For example, BatistaRoss(24) use LLMs to generate hypotheses about which headlines will lead to more user engagement (see also zhou2024hypothesisgenerationlargelanguage). Han(24)-IV prompts LLMs to generate hypotheses about candidate instrumental variables given a text description of various empirical settings. In computer science, si2024llmsgeneratenovelresearch prompt LLMs to generate new research ideas in natural language processing. [^12] Within the economics literature, discussion about how artificial intelligence more generally can be used for hypothesis generation and scientific discovery can be found in fudenberg2019predicting, ludwig2024machine, agrawal2024artificial and mullainathan2024predictive.
In this domain, the text pieces $r\in\mathcal{R}$ are the prompts for the hypothesis generation task, such as the candidate headlines or the description of the empirical setting, and $\widehat{m}(r;t)=\widehat{Y}_{r}$ is the LLM’s returned hypothesis, such as its explanation for what drives user engagement or its suggested instrumental variable. The central difference is that there is no observed economic variable $Y_{r}$ that $\widehat{Y}_{r}$ seeks to reproduce. Instead, in each hypothesis generation exercise, the researcher (sometimes implicitly) provides a scoring rule to assess the quality of the resulting hypothesis $\ell(\widehat{m}(r;t))$ and the researcher then reports the average quality of the generated hypotheses, $\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r;t))$.
Cast in this light, hypothesis generation bears striking resemblance to our analysis of the prediction problem. In assessing whether LLMs are good tools for hypothesis generation, we would like to generalize their quality in a particular research context $Q(\cdot)$ (some setting in which we would like to generate hypotheses), making inferences about $\mathbb{E}_{Q}[\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r;t)]$.
Our previous results then immediately imply that the needed guarantee is that the LLM satisfies no training leakage (Proposition 1). For hypothesis generation, no training leakage now has an interesting interpretation: do we believe that the LLM has already seen this particular prompt or setting for hypothesis generation? If so, we would naturally be concerned that the LLM is merely mimicking hypotheses it has seen before – the quality of its hypotheses on the provided task $r$ does not generalize to its performance on novel tasks. For example, in evaluating whether an LLM can generate instrumental variable strategies, we might worry that it has likely been trained on text pieces describing distance to the nearest college as a useful instrument for measuring the returns to education. In this case, observing that the LLM reproduces these existing IV strategies may not provide us with good guidance about its performance in a novel, unseen empirical setting. As before, using open source LLMs with documented training data and public weights can address the threat of no training leakage for hypothesis generation.
6.2 LLMs as Human Subject Simulators
@@
A growing body of research argues that LLMs can be used to simulate the responses of human subjects in experiments and surveys — producing what are sometimes referred to as “in-silico” subjects. For example, horton2023largelanguagemodelssimulated argues that LLMs are “implicit computational models of humans” and therefore “provide responses similar to what we might expect from a human” (pg. 1). Other studies have used LLMs to simulate human subjects in economics experiments (manning2024automatedsocialsciencelanguage; MeiEtAl(24)), in marketing (BrandEtAl(23)-LLMsForMarketResearch), in finance (Bybee(24)), in political science (Argyle_Busby_Fulda_Gubler_Rytting_Wingate_2023), and in computer science (aher2023usinglargelanguagemodels).
Using LLMs as in-silico subjects can be mapped into our framework by reinterpreting the text pieces $r\in\mathcal{R}$ and their associated economic concept $V_{r}$. Each text piece $r$ is now a particular experimental design or survey instrument and the economic concept $V_{r}$ is the average or modal response of a human subject. [^13] Of course, we could instead interpret each text piece as being associated with some random variable that summarizes the distribution of possible human subject responses on any experimental design or survey instrument. Our discussion continues to apply at the expense of more cumbersome notation. If the researcher were to collect human subject responses $V_{r}$ on a collection of experimental designs (i.e., $r\in\mathcal{R}$ such that $D_{r}=1$), she would be satisfied to calculate possible downstream parameter estimates (Equation 4).
While we could in principle collect responses on any experimental design or survey instrument (i.e., there exists some mapping $V_{r}=f^{*}(r)$ for all $r\in\mathcal{R}$), it would be costly to do so. Since LLMs could be prompted with the same experimental design or survey instrument, the researcher may instead collect the LLM’s response $\widehat{m}(r;t)=\widehat{V}_{r}$. Viewing the LLM as a computational model of human behavior, the researcher reports the plug-in parameter estimate (Equation 5) using the LLM’s responses.
To make this concrete, consider two categories of choice experiments in behavioral economics, both aiming to understand how well existing models describe the risky choices made by people.
Example: testing anomalies for theories of risky choice
In order to probe particular systematic violations of a model of risky choice, a long tradition in behavioral economics constructs “anomalies,” small collections of menus of lotteries that highlight flaws in an existing model, such as the Allais Paradox (allais1953comportement) or the original Kahneman-Tversky choice experiments (kahneman1979prospect). Researchers recruit human subjects to make choices $V_{r}$ on these specific anomalies $r$, and researchers then calculate the extent to which these choices violate, for example, expected utility theory (e.g., HarlessCamerer(94)). Could an LLM instead simulate human choices $\widehat{m}(r;t)$ on new anomalies we have yet to construct? $\blacktriangle$
Example: large-scale risky choice experiments
@@
In order to compare alternative models of risky choice, recent work encourages behavioral economists to measure the predictive accuracy of alternative models of risky choice on a diverse collection of lottery choice problems (ErevEtAl(17)-FromAnomaliesToForecasts; FudenbergEtAl(22)-Completeness). In response to this challenge, PetersonEtAl(21)-MLforChoiceTheory recruited nearly 15,000 respondents on Amazon Mechanical Turk (MTurk) to make over one million choices $V_{r}$ from menus of risky lotteries $r$, producing the “Choices13K” dataset. This impressive feat of data collection was quite costly both in terms of effort and dollars. Could an LLM instead simulate the choices of these MTurk workers $\widehat{m}(r;t)$? $\blacktriangle$
Viewing in-silico studies through the lens of our estimation framework then implies that researchers need the condition that there is no measurement error in the LLM’s outputs (Proposition 2). It implies that the LLM precisely reproduces the behavior of the target pool of human subjects on the researcher’s collection of experimental designs or survey questions.
@@
Is this assumption credible? There are studies showing LLMs can seemingly reproduce results in published experiments or surveys. However, just as with benchmark evaluations of LLMs, for every such success, there seems to be a counterexample — LLM responses to psychology experiments appear to produce more falsely significant findings than human subjects (cui2024aireplacehumansubjects), LLM responses cannot accurately reproduce the responses of human subjects on opinion polls (santurkar2023opinionslanguagemodelsreflect; boelaert_coavoux_ollion_petev_präg_2024), and LLM responses on economic reasoning tasks can be sensitive to prompt engineering (raman2024steerassessingeconomicrationality). And beyond this evidence of brittleness in the LLMs performance, there is also the second requirement for using an LLM for some estimation problem: no training leakage. The results from existing published experiments almost certainly enter into the LLMs training corpus. But presumably what we are interested in is not so much whether an LLM can memorize and regurgitate some existing result it has already seen so much as simulate human behavior in response to entirely new experiment and survey designs.
Fortunately viewing human subject simulation as an estimation problem also implies a practical fix for researchers. Just as we could collect a small validation sample to de-bias LLM labels for “standard” estimation problems (e.g., Does the ideology of a Congressional bill’s sponsor help explain the bill’s policy topic?), the same classic solution can be applied to use of LLMs to simulate human responses. One implication is that we simply cannot avoid running our experiment or survey on at least some sample of real human subjects.
Notice that the value of the LLM itself will then depend on the nature of the research application. Some economics experiments are about testing anomalies, or even a single anomaly (for example the Allais Paradox). In that case the research context $Q(\cdot)$ might contain as few as one design $r$, and the economic concept $V_{r}$ we wish the LLM to be able to reproduce is the minimum number of human subject responses we think we would need for a reliable answer to “how do people tend to choose given this menu of options $r$?” The only way we could model the LLM’s measurement error in approximating that quantity would be to collect enough human subject responses to calculate the target quantity $V_{r}$ on that very design. But then why use the LLM at all?
As a consequence, LLMs as human subject simulators are likely to only add value to experimental studies where the research context contains a large number of designs, $r$. In that setting, collecting human subject responses $V_{r}$ on a small subset of designs $r$ allows us to model the LLM’s error $\Delta_{r}$ and correct the LLM outputs on the remaining designs. For example, we could collect LLM responses on all menus in the Choices13K dataset and only collect human subject responses on a small validation sample, drastically reducing the cost of data collection. Once again, in-silico subjects can only serve to amplify, rather than fully replace, human subjects in such experimental studies.

### v3

@@ 4.3 Practical Guidance with Validation Data
Figure 3 and the top panel of Table 3 compares the average bias of the plug-in coefficient $\widehat{\beta}$ and the bias corrected coefficient $\widehat{\beta}^{debiased}$ for the target regression $\beta^{*}$ (normalized by their standard deviations) across possible combinations of bill topic $V_{r}$, linked covariate $W_{r}$, LLM $m$, and prompting strategy $p$. For most regression specifications and pairs of LLM and prompting strategy, the simple plug-in regression suffers from substantial biases. Using the validation sample yields estimates that are reliably unbiased — indeed, the bias-corrected regression coefficient performs remarkably well across all regression specifications and pairs of LLM and prompting strategy.
For each regression specification and pair of LLM and prompting strategy, Table 3 summarizes the fraction of simulations in which a 95% confidence interval centered at either the plug-in coefficient or the bias-corrected coefficient includes the target regression $\beta^{*}$. The nominal $95\%$ confidence interval for the bias-corrected regression coefficient has approximately correct coverage across all regression specifications, LLMs and prompting strategies. By contrast, plug-in estimation often suffers from severe coverage distortions.
@@
Finally, we can use these data to answer the question: If we already collected measurements $V_{r}$ in a validation sample, what is the value of the LLM labels? For each regression specification, LLM and prompting strategy, we compare the average mean square error of the bias-corrected coefficient and the validation-sample only coefficient for the target regression $\beta^{*}$. Figure 4 plots the resulting distribution across all choices of bill topic $V_{r}$, covariate $W_{r}$, and pair of model-and-prompting strategy. The average mean square error of the validation-sample-only coefficient is always higher (less precisely estimated) compared to the bias-corrected coefficient. Substantial precision improvements from using LLM outputs are possible in finite samples.
Taken together, these results indicate that estimation is a promising use case for LLMs if the researcher collects a validation sample to correct for LLM errors. LLMs can then lower the cost of data collection and improve statistical precision, while preserving the familiar econometric guarantees we desire.
4.4 Estimation without Validation Data
When using LLMs in estimation problems, validation data enables correct inference by allowing the researcher to estimate and correct for LLM errors. But what if such data cannot be collected? Before considering possible paths forward, the researcher must make a judgment call: does there exist – even in principle – a measurement procedure $f^{*}(\cdot)$ that would produce labels $V_{r}$ the researcher would trust? This is not a statistical question, but a conceptual one about the nature of the economic concept being studied.
When Ground Truth Exists
Suppose the researcher believes there exists a “true” value of the economic concept for each text piece, even if measuring it is costly or time-consuming. Consider labeling Congressional bills’ policy topics. The researcher might be confident that if they (or trained experts) carefully read each bill, they could reliably classify policy topics. The researcher may still not collect validation data despite believing ground truth exists – perhaps due to budget constraints, time pressure, or confidence in LLM accuracy. In this case, researchers might pursue one of two paths that might appear defensible but ultimately are not.
First, the researcher could argue there are no errors in the LLM’s outputs $\widehat{m}(r;t)$, justifying plug-in estimation. This argument is difficult to defend at present. A researcher proceeding this way must somehow argue why an LLM’s output exactly reproduces the economic concept, even though it is imperfect on benchmarks and why evidence on LLM brittleness does not apply.
Second, acknowledging that LLM outputs are imperfect, the researcher might write down a statistical model of its errors $\Delta_{r}=\widehat{m}(r;t)-f^{*}(r)$, just as we would in the measurement error literature. Consider a stylized model: for a given LLM $m$, the errors errors $\Delta_{r}^{m,p}=V_{r}-\widehat{V}_{r}^{m,p}$ across prompting strategies $p$ are independent. Such an assumption would suggest particular solutions: perhaps the researcher could use one prompting strategy as an instrument for another. For a given LLM $m$, its labels across prompting strategies $p$ are surely correlated with one another. Across language models $m$, there is surely substantial overlap in training datasets. Consequently, the labels $\widehat{V}_{r}^{m,p}$ are correlated (kim2025correlatederrorslargelanguage). Taking a step back, our usual measurement error frameworks were not designed for this setting – where the measurement comes from an algorithm whose behavior we do not fully understand, applied to a quantity we do not observe.
Given these challenges, the practical answer is clear: invest effort and collect a small validation sample.
When Ground Truth is Undefined
Suppose the economic concept is sufficiently abstract or subjective that the researcher is uncomfortable articulating what “ground truth” would even mean. In this case, the researcher could define the object of study as the LLM’s outputs themselves. The concept simply is, for example, whatever GPT-4o produces with a given prompt. This makes plug-in estimation valid by definition.
But the researcher is now in uncharted territory. This seemingly minor shift has important implications. When a new prompt engineering strategy is produced, should we publish a new paper? When OpenAI inevitably releases GPT-6, should we revisit all published work that used GPT-5? In Section 4.2.1, we found that variation across models and prompts can be enormous—different prompts and models can even yield estimates with different signs. If 60% of model-prompt combinations suggest a positive relationship and 40% suggest negative, what have we learned?
More fundamentally, defining the LLM’s outputs as the object of study sidesteps the hard – and economically interesting – question. We care about the underlying economic concept, not the LLM’s outputs. When facing such conceptual ambiguity, the scientific response is to acknowledge it and probe from multiple angles – not to privilege one operationalization of the concept. This is especially true with LLMs, since we have found them to be brittle and unreliable on even well-defined tasks as discussed in Section 2.3.
What might a more systematic investigation look like? As an example, suppose we compared an LLM’s outputs with measurements produced by the researcher or domain experts. How do they differ? Where do they agree and disagree? This work would help us articulate what exactly is the economic concept we are studying. By contrast, treating the LLM as ground truth bypasses this essential work: grappling with what the economic concept means and how best to measure it.
Altogether the researcher must ask themselves: Is it truly the case that no measurement procedure exists–even in principle–that they would trust? Or does one exist, but they are reluctant to invest in collecting even a small validation sample? These are fundamentally different situations. The former presents a genuine conceptual challenge deserving serious attention. The latter substitutes convenience for scientific rigor.
5 Novel Uses of Large Language Models
In addition to familiar prediction and estimation problems, LLMs enable exciting novel applications – simulating human subject responses or generating research hypotheses – that expand what we consider possible in empirical work. We offer one interpretation by mapping these applications into our framework. More broadly, these creative uses raise important questions about inferential goals, and we encourage further work to think rigorously about what researchers are ultimately trying to accomplish with such applications.
5.1 Human Subject Simulation
@@
A growing body of research uses LLMs to simulate human subjects as “in-silico” subjects across economics (e.g., horton2023large; manning2024automatedsocialsciencelanguage; MeiEtAl(24)), marketing (e.g., BrandEtAl(23)-LLMsForMarketResearch), finance (e.g., Bybee(24)), political science (e.g., Argyle_Busby_Fulda_Gubler_Rytting_Wingate_2023), and computer science (e.g., aher2023usinglargelanguagemodels).
This maps into our framework by reinterpreting text pieces $r\in\mathcal{R}$ and the existing measurement $f^{*}(\cdot)$. Each text piece $r$ represents an experimental design or survey instrument, and $V_{r}$ might represent the average or model human response. If the researcher collected human responses $V_{r}$, she would calculate downstream parameter estimates (Equation 5). But collecting human responses is costly, so instead the researcher substitutes LLM responses $\widehat{m}(r;t)=\widehat{V}_{r}$ and reports the plug-in parameter estimate (Equation 6).
Example: Testing Anomalies
To test violations of risky choice models, behavioral economists construct “anomalies” – lottery menus that highlighting flaws in the economic model, such as the Allais Paradox (allais1953comportement) or the Kahneman-Tversky choice experiments (kahneman1979prospect). Could LLMs simulate human choices $\widehat{m}(r;t)$ on new anomalies? $\blacktriangle$
Example: Large-scale Choice Experiments
@@
To compare risky choice models, recent work measures predictive accuracy across diverse lottery problems (ErevEtAl(17)-FromAnomaliesToForecasts; FudenbergEtAl(22)-Completeness). PetersonEtAl(21)-MLforChoiceTheory recruited nearly 15,000 MTurk respondents to make over one million choices $V_{r}$ from lottery menus $r$, producing the “Choices13K” dataset. Could LLMs simulate the Choices13K dataset? $\blacktriangle$
@@
Viewing human subject simulation as an estimation problem implies in-silico subjects must exhibit no measurement error (Proposition 2) — the LLM must reproduce human subject behavior on the researcher’s experiment or survey. While some studies show LLMs reproduce published experiments, counterexamples abound: LLM responses to psychology experiments appear to produce more falsely significant findings than human subjects (cui2024aireplacehumansubjects), cannot accurately reproduce the responses of human subjects on opinion polls (santurkar2023opinionslanguagemodelsreflect; boelaert_coavoux_ollion_petev_präg_2024), and can be sensitive to prompt engineering on economic reasoning tasks (raman2024steerassessingeconomicrationality). Moreover, estimation problems require no training leakage. Since published experiments likely enter LLM training corpora, the key question is whether LLMs can simulate behavior on entirely new experimental designs, not merely reproduce memorized results (manning2025generalsocialagents).
Viewing human subject simulation as an estimation problem also implies a practical fix: collect responses from at least some real human subjects. Consequently, in-silico subjects serve to amplify, rather than fully replace, human subjects. We refer the reader to recent work, such as BroskaEtAl(25); krsteski2025validsurveysimulationslimited; zhang2025agenticeconomicmodeling, which provide further guidance on debiasing in-silico subjects.
5.2 Hypothesis Generation
Recent work uses LLMs to generate hypotheses across diverse applications: predicting user engagement from headlines (BatistaRoss(24); zhou2024hypothesisgenerationlargelanguage), suggesting instrumental variables (Han(24)-IV), proposing research ideas in natural language processing (si2024llmsgeneratenovelresearch), and generating interpretable hypotheses that summarize estimated relationships between variables and text (modarressi2025causalinferenceoutcomeslearned; movva2025sparseautoencodershypothesisgeneration). We offer one interpretation viewing this as a type of prediction problem.
Researchers provide prompts $r$ containing one or more text pieces (e.g., a collection of headlines, a description of an empirical setting) and ask the LLM to generate a hypothesis $\widehat{m}(r;t)=\widehat{Y}_{r}$ that summarizes features or patterns in those texts (e.g., what drives engagement, what would be a valid instrument). The researcher evaluates average hypothesis quality $\frac{1}{N}\sum_{r\in\mathcal{R}}D_{r}\ell(\widehat{m}(r;t))$ – though this scoring rule $\ell(\cdot)$ may be implicit or informal – to determine whether the LLM is useful for hypothesis generation across different texts in the research context.
To assess whether the LLM is a useful tool for hypothesis generation in research context $Q(\cdot)$ – that is, whether the quality of its hypotheses generalizes across different texts in that context – we need no training leakage (Proposition 1). Has the LLM seen this prompt or setting before? If so, it may reproduce memorized hypotheses rather than demonstrate new capability to generate insights on novel texts. For example, an LLM trained on text describing distance-to-college as an instrument for education returns might reproduce this strategy, leading us to overestimate its ability to identify instruments in genuinely novel settings.
Hypothesis generation is an exciting frontier. Within economics, discussion about how machine learning and artificial intelligence are tools for hypothesis generation and scientific discovery can be found in fudenberg2019predicting, ludwig2024machine, agrawal2024artificial, mullainathan2024predictive, and mullainathan_rambachan_science. Further work is needed to clarify our inferential goals are in these settings – what constitutes a “good” hypothesis, and how we should evaluate LLM performance in generating them.
6 A Checklist for Empirical Research
To help use our framework in practice, we briefly summarize its key guidance for researchers incorporating LLM outputs in empirical work.
Identify Your Problem: Prediction or Estimation?
The first step is to identify your problem: are you facing a prediction problem or an estimation?
In a prediction problem (Section 3), the researcher uses text pieces $r$ to predict some linked outcome $Y_{r}$ and evaluates the LLM’s sample average loss $(1/N)\sum_{r\in R}D_{r}\ell(Y_{r},\hat{m}(r;t))$. The researcher would like to understand whether this reflects the model’s predictive performance.
In an estimation problem (Section 4), the researcher measures an economic concept expressed in text pieces $r$ to estimate downstream parameters. There is a measurement procedure $f^{*}(\cdot)$ that could be applied (e.g., the researcher reading each text piece themselves) that would produce the concept $V_{r}:=f^{*}(r)$ on all text pieces and the researcher would, for example, run a regression of $V_{r}$ on some linked covariates $W_{r}$. Since this is costly at scale, the researcher uses LLM outputs $\hat{V}_{r}:=\hat{m}(r;t)$ as a substitute.
For Prediction Problems, Ensure No Training Leakage
Valid conclusions in prediction problems require one condition: no training leakage between the LLM’s training data and the researcher’s dataset (Section 3.2).
Enforcing no training leakage is jointly determined by the prediction question and the model choice. Researchers must explicit about three key elements: (1) their target population – what collection of text pieces do they ultimately want to make predictions on? (2) their sampling procedure—how do they select evaluation samples from this population?; and (3) how does this sampling procedure relate to plausible training corpora used by the LLM? In Section 3.3, we illustrated how researchers can answer these questions in multiple common empirical settings, such as predicting on future documents, predicting on confidential documents, or constructing a random evaluation sample from a known corpus.
When in doubt, the safest approach combines open-source models with published weights or documented training data and evaluation samples constructed to mechanically exclude training data. Researchers should avoid relying on closed models like the GPT from OpenAI or the Claude from Anthropic families, since their training data is undisclosed and potentially continuously updated. Finally, prompt engineering strategies (e.g., “ignore information after date $\tau$”) are not reliable solutions to training leakage, as our evidence demonstrates.
For Estimation Problems, Collect Validation Data
In estimation problems, plugging in LLM outputs requires the strong assumption that the LLM reproduces the existing measurement procedure—an assumption that is difficult to defend given evidence on LLM brittleness in computer science (see Section 2.3 and Section 4.2).
The solution is straightforward: on a random sample of text pieces, apply the existing measurement procedure $f^{*}(\cdot)$ to collect some labels $V_{r}$. This validation sample allows the researcher to debias their downstream estimate for possible LLM errors. Section 4.3 provides a pedagogical treatment for linear regression when the LLM label appears as the dependent variable. We found that this approach delivers approximately unbiased estimates and confidence intervals with good coverage in finite samples. Moreover, the debiased estimator can be more precisely estimated and yield tighter standard errors than using the validation sample alone. In this sense, LLM outputs amplify rather than replace existing measurements.
In our simulations, validation samples with as few as 125-250 text pieces provided substantial benefits in terms of bias and coverage. The optimal size of the validation sample depends on the trade-off between the cost of collecting measurements and the potential precision gains from expanding the validation sample. This should be approached as a design choice similar to determining sample size in surveys and experiments.